(5/8)(3/8-3/16) /(1/12)
Answer by tommyt3rd(5050)

(5/8)(3/8-3/16) /(1/12)=
(5/8)(6/16-3/16) *12=
(5/8)(3/16)*12=
(5*3*12)/(8*16)=
45/32